OpenStudy (anonymous):

When solving x2 - 4x = -7 by completing the square what value is added to both sides of the equation?

OpenStudy (anonymous):

Can you tell me what is the coefficient of x here??

OpenStudy (anonymous):

you add 2 on both sides. 2ab= 2*x*2 a=x b=2

OpenStudy (anonymous):

a^2 - 2ab + b^2 = (a-b)^2